# Riccardo Giordana Pozzi

### v1

@@ 4.3 The $1/2$ BPS line defects in ${\mathcal{N}}=2$ gauge theories
Nevertheless, as discussed in subsection 3, one may still expect that some universal subsectors can be identified if, by symmetry constraints, some of the three-point functions of the individual components vanishes. Indeed, it turns out that this is the case, leaving room for universal structures to emerge.
@@
where, also in this case, the arrows represent the $Q$ actions and the Dynkin labels refers to the corresponding preserved algebra. For the displacement in SYM this is the same as 4.2, while for the ${\mathfrak{osp}}(4^{*}|2)$ one, the only difference is on the R-charge Dynkin labels associated with ${\mathfrak{su}}(2)_{{\mathcal{R}}}$.
@@
Given that the displacement enters on the three-point functions we cannot be generic on the sector to study, but we have to identify the one that do not mix. In particular, one may notice that the three-point functions
@@
The expression (4.3) makes manifest the relation between the Wilson line in ${\mathcal{N}}=2$ gauge theories and the one in SYM. This allows us to rely on the explicit evaluation of the corresponding functions of the line in SYM, [8, 14], to compute the four-point function in the present case. These are given by
@@
for the leading order and [^12] Also in this case we have rescaled by a factor of $2$, see footnote 10 and the corresponding discussion
@@
the explicit bootstrap procedure allowed to reach the expression
@@
for two unfixed parameter $c_{1}$ and $c_{2}$. Exploiting the analysis above, the universal sector allows to fix, by comparison,, the previously undetermined bootstrap parameters, leading to $c_{1}=c_{2}=-4$. With this identification, the theory can be completely solved up to NLO, as all conformal data become determined. For completeness, we report the anomalous dimensions evaluated for these fixed values of the constants
@@
Notice that, according to the argument above, this result holds universally for any line. Moreover, the analysis shows that the coefficients $c_{1}$ and $c_{2}$ are theory-independent, with the only theory-specific input in the correlator being $C_{\Phi}$. As in the case of the ${\mathcal{N}}=4$ CSm theories, the presence of unfixed parameters is simply a consequence of an insufficient number of constraints being available or employed to fully constraint the bootstrap ansatz.
We recall, however, that we have set as an initial condition the superprimary three-point function to vanish. The reason why consistency with the bootstrap result is nonetheless obtained follows from the discussion at the end of subsection 2.2. In the bootstrap procedure, the four-point function was computed by perturbing the GFF data to NLO as in (2.17) and (2.19). As discussed, under certain circumstances, this procedure automatically decouples all contributions arising from the $\mathbb{Z}_{2}$-odd sector. Indeed, in this case $\min\{\Delta\in\text{GFF}\}=2$ in (2.19), and excluding $\Delta_{\mathbb{P}}=1$ implicitly sets $\lambda_{\mathbb{P}\mathbb{P}\mathbb{P}}=0$. Given that the three-point function controls also all the $\mathbb{Z}_{2}$-odd OPEs, it follows that, effectively, the four-point function considered in the bootstrap is restricted to the $\mathbb{Z}_{2}$-even contributions, in accordance with what shown. It is possible, however, that the three-point coefficient vanishes for other reasons. We leave this, along with the study of these additional features of the correlator, for future work.
@@ 4.4 The $1/2$ BPS line in $AdS_{3}\times S^{3}\times T^{4}$
We conclude this section with a brief analysis of the cases studied in [26, 27], focusing on the comparison between the bootstrap and holographic results of [26] and the universal sector analysis. In these setups, both the second and third assumptions of Section 2 fail as one finds ${\mathcal{D}}\in{\mathcal{D}}\times{\mathcal{D}}$ as well as the presence of a tilt supermultiplet $\mathcal{T}$}, with ${\mathcal{D}}\in{\mathcal{T}}\times{\mathcal{T}}$ but ${\mathcal{T}}\notin{\mathcal{D}}\times{\mathcal{D}}$. The bootstrap analysis further shows that the displacement four-point function matches precisely the structure in (3.7). For the remainder of this discussion, however, we restrict to the simplified case where all three-point functions are set to zero, which corresponds to a dual configuration with purely Ramond–Ramond (R–R) background flux [26]. We note, however, that the full expression remains useful for further studies, as it provides a representative example of its universality class. In particular, it can be applied to other setups where universality arises, provided the operator content is consistent with that of the present case.
By setting $\lambda_{\mathbb{D}\mathbb{D}\mathbb{D}}=0$, the four-point function can be derived directly by exploiting universality as discussed in the previous examples. Concretely, we focus again on the displacement four-point function and, analogously to the analysis of subsection 4.3, we take the full trace of (4.13) in order to match the representation of the correlators.
@@
where $f^{(1)}_{\text{bulk}}(\chi)$ denotes the result obtained in [26] from the holographic computation in the case where only the R–R flux is turned on, $g$ is the bulk coupling constant and $f^{(1)}(\chi)$ is that in (4.34). By the explicit comparison one finds that
@@ 5 Universality from the holographic description
As discussed, Wilson lines can be consistently studied within the framework of supersymmetric defect CFT1, treating the line as a defect in a given SCFTd. In this section, we briefly comment on the holographic evidence supporting the analysis of the previous sections. The discussion applies only to cases that admit a string realization of the Wilson line [3, 6]. However, within this setup, the relation between different types of Wilson loops becomes more transparent and intuitive.
The discussion will largely follow the standard procedure outlined in [8, 21], while keeping the reference to the specific theory as general as possible. For clarity, however, we focus on line defects in $3d$ theories [35, 36, 37, 38], although the analysis extends straightforwardly to other dimensions, for instance, to $4d$ theories [8], leading to analogous conclusions.
@@
leaving implicit all quantities that are strictly dependent on the theories or fluxes that can be present, assuming that do not change the following considerations. Employing the standard description, we consider a Poincaré patch for the AdS4 metric
@@
where $T$ is the effective tension, $\sigma^{\mu}=(t,s)$ are the Euclidean world-sheet coordinates and we have split the embedding AdS coordinates from the compact ones. The $1/2$ BPS solution which has the Wilson line as boundary condition is achieved by considering the solution corresponding to the minimal AdS2 surface specified by
@@
where $\{w\}$ and $\{\bar{w}\}$ are generic parametrization of the metric of the compact manifold $g_{ab}(w,\bar{w})$ which can be, for example, $\mathbb{CP}^{3}$ for the ABJM case. These quantities are chosen such that they preserves a subgroup of the isometries of $\mathcal{M}$} that matches with the preserved R-symmetry expected from the dual description [^14] For example, see chapter 5 of [21]. The Nambu-Goto action can be written compactly as
@@
where $L_{2}$ is the free term, and the other are the quartic interactions, [21]. For example, $L_{4X}$ reads
@@
What is crucial for our purposes is that, although the several contributions differ from theory to theory, reflecting the dependence on the compact manifold and its associated quantities, the term $L_{4X}$ is structurally independent of the specifics of the compact manifold. In fact, evaluating the Witten diagrams for the four-point function of the transverse fluctuations $X$ and $\bar{X}$ always yields the same result, since the only relevant contribution to the connected part originates from $L_{4X}$.
From the defect CFT perspective, this implies that the displacement insertions share the same structural form irrespective of the bulk theory. Moreover, since supersymmetry relates the four-point function of the displacement to those of the other operators in the supermultiplet, the functional form of all such correlators is thereby fixed.
While the discussion above has been developed for $3d$ defects, the same reasoning applies, modulo symmetry considerations, to higher-dimensional setups, such as the $1/2$ BPS Wilson line in ${\mathcal{N}}=4$ SYM [8]. The main difference lies in the tensorial structure of the transverse fluctuations $x^{i}$ related to the different embedding of AdS2 in AdS5, instead of AdS4 as in the present case. However, the the nature of the $L_{4x}$ interaction, will be equivalent to that discussed here leading to an analysis analogous to that presented in subsection 4.1 and thus to the same conclusion.
6 The $1/2$ BPS line in ${\mathcal{N}}=2$ Chern-Simons matter theories
We now apply the approach for the identification of universal sectors, developed in section 2 and 3, to the study of $1/2$ BPS superconformal lines in $3d$ ${\mathcal{N}}=2$ theories. These cases are relevant in the context of $1/2$ BPS Wilson lines in ${\mathcal{N}}=2$ Chern–Simons matter theories [28, 38]. Moreover, since the preserved symmetry algebra in this case is a subalgebra of that preserved by the $1/6$ BPS Wilson line in ABJM [28], the two setups are closely related. In fact, the identification of universal sectors can be exploited also to gain insights into the $1/6$ BPS Wilson line at strong coupling. However, for the $1/6$ BPS line one must also take into account the additional defect supermultiplet, following an analysis similar to that outlined in section 3.
@@ 6.1 Supermultiplet
The identification of the displacement supermultiplet builds from the observation that broken generators of the bulk theory a related to local defects section by the Ward identity (2.1). In this case the $1/2$ BPS line breaks the superalgebra ${\mathfrak{osp}}(2|4)$ thus preserving half of the supercharges (A.18). For the broken ones, $\mathbf{Q}=-{}^{-}Q_{+}$ and $\bar{\mathbf{Q}}=-i{}^{+}Q_{-}$, one can associate the fermionic insertions
@@
Associated to the broken orthogonal translations, $\mathbf{P}=P_{2}+iP_{1}$ and $\bar{\mathbf{P}}=P_{2}-iP_{1}$, one instead define the displacement operators
@@
Here, the operators are labeled by $j_{0}]_{\Delta}$, where $\Delta$ is the scaling dimension and $j_{0}$ is the charge under the preserved abelian R-symmetry, (A.22). To refer to the entire multiplet, we will also use the notation of [34], indicating whether the multiplet is long $L$ ($\bar{L}$) or short at threshold AA ($\bar{A}$) with respect to $Q$ ($\bar{Q}$). The arrow represent schematically the action of the supercharge $Q$, (A.18), and explicitly this is given by
@@
Exploiting super Jacobi identities, one obtains the one involving the action of $\bar{Q}$, which would correspond to the opposite direction of the arrow in (6.1). Explicitly these relations are
@@ 6.2 Superfield realization and correlation function in superspace
such that $D\bar{y}=0$ and $\bar{D}y=0$. This allows to introduce the component expansion of generic (anti)chiral superfields obeying the conditions $\bar{D}{\mathcal{D}}(y,\theta)=0$ and $D\bar{\mathcal{D}}(\bar{y},\bar{\theta})=0$. In our case, this simply reads
@@
The coefficients has been consistently chosen by consistency with the supersymmetry action (6.2). The two point function is given by
@@
Crucially, in $1d$, although the operator insertions are ordered, it is still possible to exchange the second and fourth insertions (or, equivalently, the first and third), a process referred to as braiding. This is essentially related to a sequence of exchanging insertions, applying symmetrization, and finally using cyclicity after having identified the point at infinity, see [14]. This allows to relate different configurations, such as
@@ 6.3 OPEs and block expansions
First of all, given the correlators (6.4) and (6.5) we can see that we have different OPE to take into account and those are the chiral-antichiral ${\mathcal{D}}\times\bar{\mathcal{D}}$ and the chiral-chiral ${\mathcal{D}}\times{\mathcal{D}}$. In [28], the selection rules governing the $1/6$ BPS defect in ABJM theory were systematically derived. The analysis revealed that, in the chiral–antichiral OPE of $\mathfrak{su}(1,1|1)$, the only allowed contributions are those of the identity operator and long multiplets. Thus we have
@@
For the chiral-chiral OPE, one instead can constrained it by chirality. This simply gives
@@
This leads to a functional equivalence between the expressions for the correlators. Indeed, the combination $a_{3+n}G_{3+n}(z)$ can be mapped to $b_{3+n}G_{3+n}(\chi)$ under $z\to\chi$. This happens precisely because of the factor $(-1)^{3+n}$ in the definition of the superblock $G_{3+n}(z)$ (6.12), that compensates the corresponding sign in the parity relation (6.14). Since this argument relies only on parity, the property remains valid perturbatively, [21].

### v2

@@ 4.3 The $1/2$ BPS line defects in ${\mathcal{N}}=2$ gauge theories
Nevertheless, as discussed in Sec. 3, one may still expect that some universal subsectors can be identified if, by symmetry constraints, some of the three-point functions of the individual components vanish. Indeed, it turns out that this is the case, leaving room for universal structures to emerge.
@@
where, also in this case, the arrows represent the $Q$ actions and the Dynkin labels refer to the corresponding preserved algebra. For the displacement in SYM this is the same as in Sec. 4.2, while for the ${\mathfrak{osp}}(4^{*}|2)$ one, the only difference is on the R-charge Dynkin labels associated with ${\mathfrak{su}}(2)_{{\mathcal{R}}}$.
@@
Given that the displacement enters on the three-point functions we cannot be generic on the sector to study, but we have to identify the ones that do not mix. In particular, one may notice that the three-point functions
@@
The expression (4.3) makes manifest the relation between the Wilson line in ${\mathcal{N}}=2$ gauge theories and the one in SYM. This allows us to rely on the explicit evaluation of the corresponding functions of the line in SYM [41, 46] to compute the four-point function in the present case. These are given by
@@
for the leading order and [^13] Also in this case we have rescaled by a factor of $2$; see footnote 11 and the corresponding discussion.
@@
the explicit bootstrap procedure allowed us to reach the expression
@@
for two unfixed parameters $c_{1}$ and $c_{2}$. Exploiting the analysis above, the universal sector allows one to fix, by comparison,, the previously undetermined bootstrap parameters, leading to $c_{1}=c_{2}=-4$. With this identification, the theory can be completely solved up to NLO, as all conformal data become determined. For completeness, we report the anomalous dimensions evaluated for these fixed values of the constants
@@
Notice that, according to the argument above, this result holds universally for any line. Moreover, the analysis shows that the coefficients $c_{1}$ and $c_{2}$ are theory independent, with the only theory-specific input in the correlator being $C_{\Phi}$. As in the case of the ${\mathcal{N}}=4$ CSm theories, the presence of unfixed parameters is simply a consequence of an insufficient number of constraints being available or employed to fully constrain the bootstrap ansatz.
We recall, however, that we have set as an initial condition the superprimary three-point function to vanish. The reason why consistency with the bootstrap result is nonetheless obtained follows from the discussion at the end of Sec. 2.2. In the bootstrap procedure, the four-point function was computed by perturbing the GFF data to NLO as in (2.19) and (2.21). As discussed, under certain circumstances, this procedure automatically decouples all contributions arising from the $\mathbb{Z}_{2}$-odd sector. Indeed, in this case $\min\{\Delta\in\text{GFF}\}=2$ in (2.21), and excluding $\Delta_{\mathbb{P}}=1$ implicitly sets $\lambda_{\mathbb{P}\mathbb{P}\mathbb{P}}=0$. Given that the three-point function controls also all the $\mathbb{Z}_{2}$-odd OPEs, it follows that, effectively, the four-point function considered in the bootstrap is restricted to the $\mathbb{Z}_{2}$-even contributions, in accordance with what is shown. It is possible, however, that the three-point coefficient vanishes for other reasons. We leave this, along with the study of these additional features of the correlator, for future work.
@@ 4.4 The $1/2$ BPS line in $AdS_{3}\times S^{3}\times T^{4}$
We conclude this section with a brief analysis of the cases studied in [10, 14], focusing on the comparison between the bootstrap and holographic results of [10] and the universal sector analysis. In these setups, both the second and third assumptions of Sec. 2 fail as one finds ${\mathcal{D}}\in{\mathcal{D}}\times{\mathcal{D}}$ as well as the presence of a tilt supermultiplet $\mathcal{T}$}, with ${\mathcal{D}}\in{\mathcal{T}}\times{\mathcal{T}}$ but ${\mathcal{T}}\notin{\mathcal{D}}\times{\mathcal{D}}$. The bootstrap analysis further shows that the displacement four-point function matches precisely the structure in (3.7). For the remainder of this discussion, however, we restrict our discussion to the simplified case where all three-point functions are set to zero, which corresponds to a dual configuration with purely Ramond-Ramond (R-R) background flux [10]. We note, however, that the full expression remains useful for further studies, as it provides a representative example of its universality class. In particular, it can be applied to other setups where universality arises, provided the operator content is consistent with that of the present case.
By setting $\lambda_{\mathbb{D}\mathbb{D}\mathbb{D}}=0$, the four-point function can be derived directly by exploiting universality as discussed in the previous examples. Concretely, we focus again on the displacement four-point function and, analogously to the analysis of Sec. 4.3, we take the full trace of (4.13) in order to match the representation of the correlators.
@@
where $f^{(1)}_{\text{bulk}}(\chi)$ denotes the result obtained in [10] from the holographic computation in the case where only the R-R flux is turned on, $g$ is the bulk coupling constant and $f^{(1)}(\chi)$ is that in (4.34). By the explicit comparison one finds that
@@ 5 Universality from the holographic description
As discussed, Wilson lines can be consistently studied within the framework of supersymmetric defect CFT1, treating the line as a defect in a given SCFTd. In this section, we briefly comment on the holographic evidence supporting the analysis of the previous sections. The discussion applies only to cases that admit a string realization of the Wilson line [48, 54]. However, within these setups, the relation between different types of Wilson loops becomes more transparent and intuitive.
The discussion will largely follow the standard procedure outlined in [41, 7], while keeping the reference to the specific theory as general as possible. For clarity, however, we focus on line defects in $3d$ theories [51, 27, 6, 12, 50, 49], although the analysis extends straightforwardly to other codimensions, for instance in $4d$ theories [41], leading to analogous conclusions.
@@
leaving implicit all quantities that are strictly dependent on the theories or fluxes that can be present, assuming they do not change the following considerations. Employing the standard description, we consider a Poincaré patch for the AdS4 metric
@@
where $T$ is the effective tension, $\sigma^{\mu}=(t,s)$ are the Euclidean worldsheet coordinates and we have split the embedding AdS coordinates from the compact ones. The $1/2$ BPS solution which has the Wilson line as boundary condition is achieved by considering the solution corresponding to the minimal AdS2 surface specified by
@@
where $\{w\}$ and $\{\bar{w}\}$ are generic parametrization of the metric of the compact manifold $g_{ab}(w,\bar{w})$ which can be, for example, $\mathbb{CP}^{3}$ for the ABJM case. These quantities are chosen such that they preserves a subgroup of the isometries of $\mathcal{M}$} that match with the preserved R-symmetry expected from the dual description.  [^15] For example, see chapter 5 of [7] The Nambu-Goto action can be written compactly as
@@
where $L_{2}$ is the free term, and the other are the quartic interactions, [7]. If we focus on the four-point function of the transverse fluctuations up to NLO, only the contribution from $L_{4X}$ is needed and explicitly it reads
@@
Crucially, $L_{4X}$ is structurally independent of the specifics of the compact manifold. Therefore, the evaluation of the Witten diagrams for the four-point functions of the transverse fluctuations $X$ and $\bar{X}$ always yields the same result, as the only relevant contribution to the connected part comes from $L_{4X}$.
From the defect CFT perspective, this implies that the displacement insertions share the same structural form irrespective of the bulk theory. Moreover, since supersymmetry relates the four-point function of the displacement to those of the other operators in the supermultiplet, the functional form of all such correlators is thereby constrained, up to NLO, and different correlation functions can be related as in the previous sections; see Sec. 4.1.
While the discussion above has been developed for defects in $3d$ theories, the same reasoning applies, modulo symmetry considerations, to different line defects in higher-dimensional setups, such as the $1/2$ BPS Wilson line in ${\mathcal{N}}=4$ SYM [41]. The main difference lies in the tensorial structure of the transverse fluctuations $x^{i}$ related to the different embedding of AdS2 in AdS5, instead of AdS4 as for the cases discussed above. However, the nature of the $L_{4x}$ interaction, discussed in [41], will be qualitatively equivalent to that discussed here. Therefore, the four-point interactions can be identified leading to an analysis analogous to that presented in Sec. 4.2.
6 The $1/2$ BPS line in ${\mathcal{N}}=2$ Chern-Simons-matter theories
We now apply the approach for the identification of universal sectors, developed in Secs. 2 and 3, to the study of $1/2$ BPS superconformal lines in $3d$ ${\mathcal{N}}=2$ theories. These cases are relevant in the context of $1/2$ BPS Wilson lines in ${\mathcal{N}}=2$ Chern-Simons-matter theories [8, 49]. Moreover, since the preserved symmetry algebra in this case is a subalgebra of that preserved by the $1/6$ BPS Wilson line in ABJM [8], the two setups are closely related. In fact, the identification of universal sectors can be exploited also to gain insights into the $1/6$ BPS Wilson line at strong coupling. However, for the $1/6$ BPS line one must also take into account the additional defect supermultiplet, following an analysis similar to that outlined in Sec. 3.
@@ 6.1 Supermultiplet
The identification of the displacement supermultiplet builds from the observation that broken generators of the bulk theory are related to local defects section by the Ward identity (2.1). In this case the $1/2$ BPS line breaks the superalgebra ${\mathfrak{osp}}(2|4)$ thus preserving half of the supercharges (A.18). For the broken ones, $\mathbf{Q}=-{}^{-}Q_{+}$ and $\bar{\mathbf{Q}}=-i{}^{+}Q_{-}$, one can associate the fermionic insertions
@@
Associated to the broken orthogonal translations, $\mathbf{P}=P_{2}+iP_{1}$ and $\bar{\mathbf{P}}=P_{2}-iP_{1}$, one instead defines the displacement operators
@@
Here, the operators are labeled by $j_{0}]_{\Delta}$, where $\Delta$ is the scaling dimension and $j_{0}$ is the charge under the preserved Abelian R-symmetry, (A.22). To refer to the entire multiplet, we will also use the notation of [2], indicating whether the multiplet is long $L$ ($\bar{L}$) or short at threshold AA ($\bar{A}$) with respect to $Q$ ($\bar{Q}$). The arrow represents schematically the action of the supercharge $Q$, (A.18), and explicitly this is given by
@@
Exploiting super-Jacobi identities, one obtains the one involving the action of $\bar{Q}$, which would correspond to the opposite direction of the arrow in (6.1). Explicitly these relations are
@@ 6.2 Superfield realization and correlation function in superspace
such that $D\bar{y}=0$ and $\bar{D}y=0$. This allows us to introduce the component expansion of generic (anti)chiral superfields obeying the conditions $\bar{D}{\mathcal{D}}(y,\theta)=0$ and $D\bar{\mathcal{D}}(\bar{y},\bar{\theta})=0$. In our case, this simply reads
@@
The coefficients have been consistently chosen by consistency with the supersymmetry action (6.2). The two-point function is given by
@@
Crucially, in $1d$, although the operator insertions are ordered, it is still possible to exchange the second and fourth insertions (or, equivalently, the first and third), a process referred to as braiding. This is essentially related to a sequence of exchanging insertions, applying symmetrization, and finally using cyclicity after having identified the point at infinity, see [46]. This allows one to relate different configurations, such as
@@ 6.3 OPEs and block expansions
First of all, given the correlators (6.4) and (6.5) we can see that we have different OPEs to take into account and those are the chiral-antichiral ${\mathcal{D}}\times\bar{\mathcal{D}}$ and the chiral-chiral ${\mathcal{D}}\times{\mathcal{D}}$. In [8], the selection rules governing the $1/6$ BPS defect in ABJM theory were systematically derived. The analysis revealed that, in the chiral-antichiral OPE of $\mathfrak{su}(1,1|1)$, the only allowed contributions are those of the identity operator and long multiplets. Thus we have
@@
For the chiral-chiral OPE, one instead can constrain it by chirality. This simply gives
@@
This leads to a functional equivalence between the expressions for the correlators. Indeed, the combination $a_{3+n}G_{3+n}(z)$ can be mapped to $b_{3+n}G_{3+n}(\chi)$ under $z\to\chi$. This happens precisely because of the factor $(-1)^{3+n}$ in the definition of the superblock $G_{3+n}(z)$ (6.12), that compensates the corresponding sign in the parity relation (6.14). Since this argument relies only on parity, the property remains valid perturbatively [7].